Maria bicycles 5 km/h faster than Carlos. in the same time it takes Carlos to bicycle 6 km Maria can bicycle 7 km. how fast does each bicyclist travel?
Answer by TimothyLamb(4379)

x = maria speed
y = carlos speed
---
x = y + 5
---
s = d/t
t = d/s
---
maria time:
t = 7/x
t = 7/(y + 5)
---
carlos time:
t = 6/y
---
equate times:
7/(y + 5) = 6/y
7y = 6(y + 5)
7y = 6y + 30
y = 30
x = 35
---
answer:
x = maria speed = 35 kph
y = carlos speed = 30 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php